Skip to main content

AAVSOnet status for 2014-07-02

HQA's picture
HQA
Offline
Joined: 2010-05-10

The computer breach stopped our processing for several weeks, as we needed to make the connections between telescope site and HQ more secure.  As of now, most of the images taken in June have been transferred to HQ, and we're working on the processing.  Anything after June 4 is not yet available, though the image acquisition continues unabated.

BSM_HQ.  We have the system configured, but are in the midst of doing polar alignment with the new Celestron AVX mount that was donated by Richard Berry.  We hope to have this system on-line next week.  Testing the Aurora cloud monitor.

BSM_Argentina.  Waiting for the delivery of a new hand paddle.

BSM_South.  Working as normal.

BSM_Berry.  Shipped to Perth.  Upon arrival, some damage was noted.  Greg Bolt thinks that he can make repairs, so that the system will be functional.  Probably another month or so before the enclosure is finished and commissioning can begin.

BSM_Hamren.  Had a failure of an RA drive plastic gear.  Celestron is sending a replacement.  Mike Linnolt will continue testing after that gear is installed.  We are planning a move to his better (drier) remote site in a few months.

Coker30.  System running fine; will be off-line for a few days over the 4th.

SRO.  System running fine; monsoon shutdown expected in a week or so.

TMO61.  Shut down for monsoon.  We hope to use this period to refurbish the top end so that the LISA spectrograph donated by Bart Staels can be installed.  There may be a rare night or two, but expect the system to be off-line until September.

OC61.  Working fine.  Nigel Frost is working on the instrument selector; all parts for the eShel spectrograph are at the observatory.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484